Given: 3x - 8 = -1 and -5x + 8y = 7, solve for x and y.
Answer by oberobic(2304)

I assume you left the 'y' out of the first equation.
.
3x -8y = -1
-5x +8y = 7
------------ add
-2x = 6
x = -3
.
3(-3) -8y = -1
-8y = 8
y = -1
.